OpenStudy (anonymous):

A wheel 90.0 cm in diameter accelerates uniformly from 205 rpm to 390 rpm. If in this process a point on the edge of this wheel have traveled 106 m, how long does this process take? (seconds)

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

\[d= r \theta \]

OpenStudy (anonymous):

\[\frac{ 106 }{ .9} = \theta \]

OpenStudy (anonymous):

convert theta to rev.s and use: \[\omega _{f}^{2} =\omega _{i}^{2} +2\alpha \theta \] to find alpha

OpenStudy (anonymous):

then find time in minutes from: \[\omega _{f} = \omega _{i} + \alpha t\]